Help Save &Exit Submlt How many years (and months) intermediate calculations. Round months to 1 decimal place.) will it take $2 million to grow to $6.20 million with an annual interest rate of 7 percent? (Do not round
Answer #1

Amount invested = $2 million = $2,000,000

Future value = $6.20 million = $6,200,000

Annual interest rate = r = 7% = 0.07

Let the number of years = n

Future value = Amount invested * (1+r)n

$6,200,000 = $2,000,000*1.07n

1.07n = $6,200,000/$2,000,000 = 3.10

n*log 1.07 = log 3.10

n = log 3.10/log 1.07

n= 0.491362/0.029384 = 16.72

Number of months after 16 years = 12*0.72 = 8.65 = 9 months

Hence, answer is 16 years and 9 months